# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

### 05 Feb 2024 v1

@@ 5.2. Analysis
To measure the effect of the ReviewFlow system on each dimension of the review quality (eg. summative, coverage, justified), we conducted repeated measure ANCOVA tests. We used paper id, the order of experiment conditions (whether ReviewFlow was used for the participant’s first or second paper), the knowledge level of each paper topic, and the length of each review as co-variants.
To measure the effects of the experiment condition on the time they spent reading independent reviews and writing reviews taking the consideration of the differences between the paper topic and order effects, we ran a repeated measure ANCOVA using the paper id, the order of experiment condition and the knowledge level of each paper topic as co-variants.
@@
All semi-structured interviews with participants are recorded and transcribed. Two researchers conducted iterative open coding on the transcripts using Dovetail [^4] https://dovetailapp.com/ and conducted thematic analysis on the transcripts (Braun and Clarke, 2006). They open-coded the data by identifying topics mentioned by the participants. Initial codes were combined into preliminary themes, which were discussed among the research team. After iteratively discussing the code themes, researchers derived the final themes around: participants’ reactions to each feature, their overall perceptions of the ReviewFlow system, and their concerns about using the system.
@@ 6. Results
We report our results from the within-subjects experiment and the interview study. In the within-subject experiments, across both conditions, participants spent an average of 38 minutes writing a review with an average length of 243 words written. 59.3% of participants chose to accept decisions, which is consistent with the original decisions for the two papers. Our findings suggested that ReviewFlow provides more guidance to novice reviewers and can make the review process more useful and engaging.
6.1. RQ1: ReviewFlow helps improve the review draft quality
6.1.1. ReviewFlow helps participants write more comprehensive reviews, as rated by experts
To compare the quality of written review in both conditions, we performed repeated measures ANCOVA to examine the effect of the two conditions on the review length as well as each quality measure. We control the order of experiment condition and the knowledge level of each paper topic as co-variates. As shown in Figure 7, we found that the written reviews in the ReviewFlow condition are statistically significantly more comprehensive reviews than the baseline condition(F = 3.55, p = 0.04 *** ). We found no significant interaction effect between the order of the experiment conditions and no statistically significant differences between the knowledge of the two papers.
Figure 7. Experts evaluation on the quality of review written by ReviewFlow versus Baseline.
We also found that participants wrote longer reviews and called out more strengths and weaknesses in the paper in the ReviewFlow condition. However, there is no significant improvement in the constructiveness of the review. This indicates that the ReviewFlow system can scaffold participants to capture more pros and cons but still cannot help participants write constructive items towards each weakness. In addition, we asked the participants to self-rate their satisfaction with the review quality. Participants are more satisfied with the reviews written in the ReviewFlow condition.
ReviewFlow Baseline
@@ 6.1.1. ReviewFlow helps participants write more comprehensive reviews, as rated by experts
Num of strengths
2.38(0.19) 1.91 (0.17)
Num of weaknesses
2.62(0.22)) 2.08(0.27)
Length(words)
250.8(19.4) 235.8(23.5)
Self-rated satisfaction on the review
4.35(0.4) 4.0(0.34)
Time-on-task
reading time (minutes)
26.9 (1.9) 19.2(2.1)
writing time (minutes)
15.5(2.2) 15.7(2.4)
Table 1. Proxy of written review quality includes number of strengths, number of weaknesses, and length of the review. Participants self-perceived review quality. Time on tasks in both conditions
6.2. RQ2: Longer interaction duration with ReviewFlow but improves users’ self-efficacy
@@
We performed repeated measures of ANCOVA to examine the effect of the two conditions on writing time and control the length of the review as a co-variate. The results did not show a significant difference. We observed that ReviewFlow reduced review writing time but the reading time is higher compared to the baseline, as shown in Table 1. However, even though participants spent more time on reading, they reported that “ReviewFlow saved me more time on writing review”. One participant reflected on the reasons that “I feel actively receiving these pop-up questions make me spend more time on reading the paper, judging it from multiple dimensions and taking notes, but I think it did save me time on the review drafting since I don’t need to go through all the notes again”(P2).
@@ 6.2.2. Participants reported to have higher self-efficacy after experiencing ReviewFlow than the Baseline
To compare self-efficact in both condition, we conducted the Mann- Whitney U Test between each participant’s ratings on self-efficacy given the sample is small (McKnight and Najab, 2010). Results showed that self-efficacy ratings in the ReviewFlow (M = 4.92, STD = 0.40) is significantly higher than the Baseline (M = 3.92, STD= 0.36, p = 0.05 ***)
6.2.3. Participants’ took more notes in the ReviewFlow and used the features actively
Participants took significantly more notes in the ReviewFlow system (M=12.9, STD = 3.7) than baseline environment (M = 9.1, STD = 3.9, Wilcoxon signed-rank test, Z= 103.0, p ¡= 0.01, ***** *). This indicates that participants are more engaged in the reading process with ReviewFlow and not only reading text but also critically reflecting on the content. All participants used the notes-to-outline features which helped them summarize notes into an outline. Three participants did not expand the high-level outline to a detailed outline.
# of participants Avg frequency useful accuracy
Used the section-specific reflection question 14 3.8 5.4 4.9
Used the contextual reflection question on highlight 10 1.6 3.8 4.3
Checked the in-situ citation support 10 1.0 4.6 6.2
Checked the missing citation support 9 1.0 2.5 3.4
Used the summarization notes to high-level outline 16 1.12 6.3 5.2
Expanded the high-level outline to detailed outline 13 1 5.2 4.3
Table 2. Usage of scaffolding features and participants’ rating on the usefulness and accuracy. Average frequency is the average number of times the feature was used. The notes-to-outline feature is used by all participants, while some participants did not expand the high-level outline to a detailed outline.
@@ 6.3. RQ3: Participants perceived ReviewFlow as useful in the review workflow
After participants had experienced the two systems, we asked participants to compare them overall directly. As shown in Figure 8, participants highly prefer ReviewFlow over the baseline environment. 92.9% of participants perceived the interface as enjoyable to use. Participants mentioned that the scaffolding features make the review process more engaging and less boring (N=5).
Figure 8. Participants’ reactions to the two systems in terms of easiness, enjoyment, and efficiency
Figure 9. Participants’ perceptions of the engagement, cognitive workload and feeling of control in two conditions
We further measured participants’ perceptions of conducting the task together with the system. In the post-survey, we asked participants to rate cognitive workload, such as distraction and engagement, the feeling of control and collaboration with the interface, and perceived learning outcomes with the interface, as shown in Figure 9. The add-on scaffolded features bring in some distraction since 57.1% of the participants believe that the interaction with the interface distracts them from the review. Interestingly, participants reflected that “It is actually not a bad distraction, since you don’t want to keep reading the paper and get nothing out of it. This distraction is like a pop-up that keeps telling you about review criteria from different places”(P8). Even though the current system uses AI-generated features, such as notes-to-outline synthesis, all participants agree that they still have the control over writing process. One participant mentioned the reason is that “It just synthesizes what I have already written, and if I want, I can just delete them”(P9).
64.3% of participants have the feeling of collaborating with the interface. Participants identified the collaboration mainly happened in the process of answering reflections on each section by asking and answering questions and note-to-outline synthesis. The two-step process where participants can first summarize notes into a high-level outline and then edit and expand it provides participants the feeling of “iterating with an assistant”(P8). When we asked participants their reason for being biased by the outputs, most participants thought that the notes synthesis did not bias them in the current session, but they worried that “lazy people might randomly create notes and use the outline to make a decision ”(P6). In addition, all participants believe that the ReviewFlow interface helped them learn the peer review process.
@@ 6.3.1. Participants have different preferences on the scaffolding strategies
As shown in Table 2, participants perceived that section-specific reflections were more useful and more accurate than the contextual question on the highlighted text. Participants reflected (N=2) that the uncertainty on the contextual questions is high since the questions’ quality depends on how much text they highlight. P15 also reflected that “I kind of have a question in my mind when I highlight certain parts of the text, so if the question does not match with the question I was thinking, I feel this is not useful”. We also found that many people did not pay attention to the missing citation support, since they did not focus on evaluating the related work section while reading. The most useful feature perceived by participants is the notes-to-outline synthesis feature where participants reflected that it is “very easy to get a sense on top of my notes”(P3). Participants reflected that the detailed outline bullets contain the same meaning as the high-level outline topic which makes it less useful. We further conducted an in-depth analysis of how participants used and perceived different types of scaffolding strategies:
Contextual reflections cues worked as an expert to guide reviews evaluated from different criteria
Participants described that their process of using the section-specific reflections is slightly different from the process of using context-specific reflections. P11 described that “I first check out the questions real quick before each section, and then I dive into the reading for this section. After I finish reading it, I come back and revisit these questions and reflect on how I feel”. While, for contextual questions, “it is more like I feel confused or not sure what I should think about, I will highlight and use those questions to guide me”(P11). P14 described the section-specific question as “an expert who is sitting there that keeps giving you guided questions and prompts you to think”(P14).
@@
All participants used the notes-to-outline synthesis. Participants reflected that they prefer to have the notes in a “structured version” and that they don’t need to conduct the mapping by hand (N=5). However, there are three participants did not use the expand button to obtain the detailed outline. Two participants intentionally avoided using it since they were worried the detailed bullets generated by AI may mislead them in their decision-making.
6.3.2. Participants’ potential concerns of using ReviewFlow
When we asked participants about their main concerns about using ReviewFlow, they mostly concentrated on the potential errors that AI can make in the notes synthesis process. Participants did not express severe concerns in the study session since the generated outline is mainly reusing or summarized their notes. However, they were still worried that the nuanced tone in the outline produced by AI may exaggerate the weakness of the paper and influence reviewers’ perceptions. For instance, one participant mentioned that if you wrote the notes as “the details around study participants seems a little bit unclear”, but the AI turns it into “this paper lack clarity”, which may be misleading to the reviewer. Moreover, participants expressed their concerns of other reviewers in the community. They are worried that “lazy reviewers or last minute reviewers may randomly leave notes” and then use the tool to generate an outline. This misuse of the tool will not be fair to the paper authors and the research community in general.
@@ 7. Discussion
Advances in large language models are changing how work gets done. Our research explores how we might integrate intelligent scaffolding in a way that informs and guides novices and steers away from biasing the underlying judgement. To inform the design of intelligent scaffolding, we conducted a formative study where we learned what strategies and practices experts adopt, as well as, what challenges novices face when writing peer reviews. Our ReviewFlow system aimed to leverage LLMs to provide timely considerations while reading and assessing a submission and structural support when composing a review. Our within-subject experiment (N=16 participants) found that novice reviewers not only preferred ReviewFlow over the baseline system, but they also wrote longer and more comprehensive reviews, as judged by experts. Using ReviewFlow participants spent more time reading the paper and took more detailed notes. Those notes were aided by ReviewFlow’s contextual cues, and helped participants call out more strengths and weaknesses. Using ReviewFlow, participants were more satisfied with their reviews (according to self-ratings) and attributed this to the timely cues and a useful workflow.
In comparison with general peer review practices (Wang et al., 2020; Shah, 2022), participants using ReviewFlow allocated more time to reflect on each section and evaluating the paper’s quality. Interestingly, 40% of participants opted for rejection decisions in both conditions, contrary to the original decisions for the two papers. A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions (Moore, 2013). However, participants in both conditions still found it hard to provide constructive feedback on each weakness, even with example reviews. Perhaps future systems could incorporate additional features into the intelligence scaffolding, such as presenting examples of improvement suggestions from similar prior reviews or offering feedback to the reviewers after they have an initial draft (Krause et al., 2017).

### 26 Feb 2024 v2

@@ 5.2. Analysis
To measure the effect of the ReviewFlow system on each dimension of the review quality (eg. tone, comprehensive, constructive and justified), we conducted repeated measure ANCOVA tests. We used paper ID, the order of experiment conditions (whether ReviewFlow was used for the participant’s first or second paper), the knowledge level of each paper topic as co-variants. To measure the effects of the experiment condition on the time they spent reading independent reviews and writing reviews taking the consideration the differences between the paper topic and order effects, we ran a repeated measure ANCOVA using the paper ID, the order of experiment condition and the knowledge level of each paper topic and the review word counts as co-variants. To compare participants’ behaviors between the two conditions, we conducted Wilcoxon signed-rank test, which is a non-parametric rank test, on users’ interaction data, e.g. number of notes participant taken in both conditions (Woolson, 2007).
@@
All semi-structured interviews with participants are recorded and transcribed. Two researchers conducted iterative open coding on the transcripts using Dovetail [^6] https://dovetailapp.com/ and conducted thematic analysis on the transcripts (Braun and Clarke, 2006). They open-coded the data by identifying topics mentioned by the participants. Initial codes were combined into preliminary themes, which were discussed among the research team. After iteratively discussing the code themes, researchers identified the final themes around: participants’ reactions to each feature, their overall perceptions of the ReviewFlow system and their reactions to each scaffolding feature.
@@ 6. Results
We report our results from the within-subjects experiment and the post-interview. In the within-subject experiment, across both conditions, participants spent an average of 38.7 minutes writing a review with an average length of 243 words written. 60% of the participants decided to accept the paper, which is consistent with the original decisions for the two papers. Our findings revealed that ReviewFlow provided more guidance to novice reviewers and made the review process more useful and engaging.
6.1. RQ1: ReviewFlow helped participants write more comprehensive reviews
To compare the quality of written review in both conditions, we performed ANCOVA tests to examine the effect of the two conditions on each quality measure, accounting for the review length, participants’ knowledge of the topic, the order of the experiment condition as co-variants. As shown in Figure 7, we found that the reviews in the ReviewFlow condition are significantly more comprehensive than the Baseline condition (p = 0.04***, F = 3.55). We found no significant interaction effect between the order of the experiment conditions and no statistically significant interaction effect between the knowledge of the two papers.
Figure 7. Experts’ evaluation scores on the quality of review written by ReviewFlow versus Baseline.
Furthermore, we evaluated several proxies of review quality including the number of strengths and weaknesses and participants’ self-rated satisfaction with the written review. As shown in Table 1, participants wrote longer reviews and called out more strengths and weaknesses in the paper in the ReviewFlow condition. However, there is no significant improvement in the constructiveness of the review. This indicates that the ReviewFlow system can scaffold participants to capture more pros and cons but still cannot help participants write constructive solutions for each weakness. Also, participants’ self-rated satisfaction with reviews written using the ReviewFlow tends to be higher than using the Baseline. ANCOVA test with proxies did not a show significant effect. Participants reflected on the reason that “I captured more pros and cons since I took more notes this time, so I feel satisfied with the review as it covered more aspects” [P3].
ReviewFlow Baseline p F
@@ 6.1. RQ1: ReviewFlow helped participants write more comprehensive reviews
Count of strengths 2.38 (0.19) 1.91 (0.17) - -
Count of weaknesses 2.62 (0.22)) 2.08 (0.27) - -
Self-rated satisfaction 4.35 (0.4) 4.00 (0.34) - -
Length of the review
Word counts 250.8 (19.4) 235.8 (23.5) - -
Time on task
Reading time (minutes) 26.9 (1.9) 19.2 (2.1) *** 19.2
Writing time (minutes) 15.5 (2.2) 15.7 (2.4) - -
Self-efficacy
Self-efficacy on reviewing 4.92 (1.43) 3.92 (1.36) * 4.2
Table 1. Proxies of review quality include the number of strengths, number of weaknesses, and satisfaction on the written review rated by participants themselves. Participants included more weaknesses in the ReviewFlow condition. Participants spent more time reading and reported higher self-efficacy after using the ReviewFlow.
6.2. RQ2: Longer interaction duration with ReviewFlow but improved participants’ self-efficacy
@@
We performed ANCOVA to examine the effect of the two conditions on writing time, accounting for the length of the review, participants’ knowledge of the topic, and the order of the experiment condition as covariates. As shown in Table 1, participants spent significantly more time in reading and sense-making with the ReviewFlow than with the Baseline. However, they spent less time on writing, while the result was not significant. 68.8% of them still reported that “ReviewFlow saved me more time on writing reviews”. P11 reflected the reasons that “Answering pop-up questions made me spend more time on reading the paper, judging it from different aspects, and taking notes, but I feel it did save me time in the end since I don’t need to go through all the notes again”[P11]. Similarly, P6 also mentioned the reason for saving time as “having that[my notes] summarized and organized at the end meant that I didn’t have to go back through each section and think about where I should put the strengths and weaknesses”[P6]. On the contrary, P5 reflected that they spent more time considering different aspects, such as validity, novelty, clarity, etc. P5 highlighted that “I can write the review fast, but that is not my goal. As a reviewer, it is more important to spend enough time carefully evaluating the paper”[P5].
@@ 6.2.2. Participants reported to have higher self-efficacy after experiencing ReviewFlow than the Baseline
We asked participants to rate their self-efficacy on the ability to conduct a conference peer review and conducted the ANCOVA test between each participant’s ratings, accounting for covariates. Table 1 showed that self-efficacy ratings in the ReviewFlow are significantly higher than the Baseline. This result indicated that participants built more confidence in learning after using the ReviewFlow. Specifically, participants described their learning process –“I feel I gradually built confidence. At the beginning, I would try to answer every question. But later on, I started to remember which aspect I should think about while reading that section, so I didn’t need to frequently check guided questions”[P13].
6.2.3. Participants took more notes in the ReviewFlow and used the features actively
Participants took significantly more notes in the ReviewFlow condition (M=12.9, STD = 3.7) than in the Baseline condition (M = 9.1, STD = 3.9, Wilcoxon signed-rank test, Z= 103.0, p ¡= 0.01, ***** *). This indicates that participants are more engaged in the reading process with ReviewFlow and not only reading text but also critically reflecting on the content. Participants actively answered the guided questions. All participants used the notes-to-outline features which helped them summarize these notes into an outline. Three participants did not expand the high-level outline to a detailed outline.
@@ 6.3. RQ3: Participants perceived ReviewFlow as useful in the review workflow
After participants had experienced the two systems, we asked them to compare two conditions. As shown in Figure 8, participants highly preferred ReviewFlow over the Baseline. 93.7% of participants perceived the interface as enjoyable to use. Participants mentioned that the scaffolding features make the review process more engaging and less boring (N=5). All participants think that the system is useful and they like the guidance provided in reviewing. For instance, P11 described the guidance they received as similar to experts – “I feel like some experts, such as my advisor, were sitting next to me to prompt me in particular ways and show me how to write the review”[P11].
Figure 8. Participants’ reactions to the system in two conditions in terms of easiness, enjoyment, and efficiency.
We further measured participants’ perceptions of the tasks in each condition. In the post-survey, we asked participants to rate cognitive workload, such as distraction and engagement, the feeling of control and collaboration with the interface, and perceived learning gains. As shown in Figure 9, 75% of the participants reported that they are engaged in the process, but the add-on scaffolding features bring in some distraction for 68.8% of the participants. However, participants highlighted that it is not a bad distraction, but served as a staging process that motivated them to think. For instance, P8 explained that “Since you don’t want to just keep going through the paper without getting anything from it. The distraction is like a pop-up that keeps prompting you about review criteria from different places”[P8].
All participants believe that the ReviewFlow helped them learn the peer review process. 68.8% of participants have the feeling of collaborating with the interface. Participants identified the collaboration mainly happened during the process of answering contextual cues in each section and using the note-to-outline synthesis feature. The two-step process where participants can first summarize notes into a high-level outline and then expand it provided them the feeling of “iterating with an assistant”[P9]. Even though the current system uses AI-generated outlines, all participants agreed that they still have the control over writing process. One participant mentioned the reason is that “It just synthesized what I have already written in the notes and I can still write it by myself”[P9]. 37.5% of the participants described that the notes synthesis did not bias them in the current session, but they still worried that “busy reviewers might randomly create notes and use the outline to make a decision”[P14].
Figure 9. Participants’ feelings of the engagement, control, and perceptions of bias in two conditions.
@@ 6.3.1. Participants have different preferences on the scaffolding strategies
As shown in Table 2, participants perceived that section-level cues were more useful and more accurate than the phrase-level cues on the highlighted text. Two participants reflected that the uncertainty on the phrase-level cues is high since the questions’ quality depends on the number of words they highlighted. P15 also said “I have already had a question in my mind when I highlighted certain parts of the text. So if the question did not match with the question I was thinking, I felt a bit disappointed”[P15]. We also found that many people did not pay attention to the missing citation support, since they did not focus on evaluating the related work section. The most useful feature perceived by participants is the notes-to-outline synthesis feature where participants reflected that it helped them “easily get a sense on top of my notes”[P3]. However, three participants did not expand the topic to a detailed outline. They reflected that the detailed outline bullets contained similar content as the high-level outline topic which made it less useful. We further conducted an in-depth analysis of how participants used and perceived different types of scaffolding strategies:
#(p) freq useful accuracy
Section-level cues
14 3.8 5.4 4.9
phrase-level cues
10 1.6 3.8 4.3
In-situ citation pop-up
10 1.0 4.6 6.2
Missing citation pop-up
9 1.0 2.5 3.4
Summarize notes to high-level topics
16 1.12 6.3 5.2
Expanded topics to detailed outline
13 1 5.2 4.3
Table 2. Usage of scaffolding features shows the number of participants, average frequency, and participants’ rating on the usefulness and accuracy. All participants used the features that synthesize notes to high-level topics, while three participants did not expand it to a detailed outline.
Contextual cues guided novices to evaluate from different criteria
Participants described that the contextual questions are “more specific and tailored to the paper compared with the general community instruction”(N=3). Instead of providing the review criteria with general questions, such as “how well does the paper execute their contribution?”, participants preferred to have more contextually guided questions according to the paper content, such as “does the method section provide clear and well-justified explanations of the research prototype’s design including the choice of videos and the method to set up the activity recognizer?”. Participants described that their process of using the section-level cues is slightly different from the phrase-level cues. P10 described that “I quickly checked out the questions next to each section before I dived into the reading. After I finished reading the section, I came back to these questions and reflected on how I felt”[P10]. Correspondingly, phrase-level questions are used when “I feel confused or not sure what I should think about”[P11]. These contextual cues reminded them or prompted them to evaluate different aspects of each section.
@@
All participants used the notes-to-outline synthesis. Participants preferred having notes in a “structured version” and indicated that they didn’t need to summarize and map their notes by hand. The ability to summarize notes into each review section helped them ”reflect on the strengths and weaknesses” (N=2). As a result, they perceived that the outline generation feature saved their time in writing (N=5). Participants liked the fact-checking function that when they can map each topic back to their notes and original pdf, as shown in Figure 6. The visual mapping helped them cross-check the notes and increased their trust in the topic. However, three participants intentionally avoided using the expand button to get a more detailed bullet. They are concerned that the expanded outline may generate content from nowhere. P3 explained that “I was worried that if I clicked to expand the themes into an outline, it would have incorporated things that I didn’t want to include”[P3].
6.3.2. Participants’ concerns about using ReviewFlow
Participants are mostly concerned about the potential errors that AI can make in the notes synthesis process. Participants did not express severe concerns in the study session since the generated outline mainly reused or summarized their notes. However, they were still worried that the nuanced tone in the outline produced by AI may exaggerate the weakness of the paper and influence reviewers’ perceptions. For instance, one participant mentioned that if you wrote the notes as “the details around study participants seems a little bit unclear”, but the AI turns it into “this paper lack clarity”, which may mislead the reviewer. Moreover, participants expressed their concerns of other reviewers in the community. They are worried that “last minute reviewers may randomly leave notes and then use the tool to generate an outline. If they expand the notes it might not be truly reflective of what they thought if they were not paying attention. The misuse of the tool will not be fair to the paper authors and the research community in general”[P14].
@@ 7. Discussion
Advances in large language models are changing how work gets done. Our research explores how we might integrate intelligent scaffolding in a way that informs and guides novices and steers away from biasing the underlying judgment. To inform the design of intelligent scaffolding, we conducted a formative study where we learned what strategies and practices experts adopt, as well as, what challenges novices face when writing peer reviews. Our ReviewFlow system aimed to leverage LLMs to provide timely considerations while reading and assessing a submission and structural support when composing a review. Our within-subject experiment (N=16 participants) found that novice reviewers not only preferred ReviewFlow over the baseline system, but they also wrote longer and more comprehensive reviews, as judged by experts. Using ReviewFlow participants spent more time reading the paper and took more detailed notes. Those notes were aided by ReviewFlow’s contextual cues, and helped participants call out more strengths and weaknesses. Using ReviewFlow, participants were more satisfied with their reviews (according to self-ratings) and attributed this to the timely cues and a useful workflow.
In comparison with general peer review practices (Wang et al., 2020; Shah, 2022), participants using ReviewFlow allocated more time to reflect on each section and evaluate the paper’s quality. Interestingly, 40% of participants opted for rejection decisions in both conditions, contrary to the original decisions for the two papers. A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions (Moore, 2013). However, participants in both conditions still found it hard to provide constructive feedback on each weakness, even with example reviews. Perhaps future systems could incorporate additional features into the intelligence scaffolding, such as presenting examples of improvement suggestions from similar prior reviews or offering feedback to the reviewers after they have an initial draft (Krause et al., 2017).